substitution method to solve equation
x+4y=5
2x+13y=15

Answer by robertb(5830)

x + 4y = 5 ==> x = 5 - 4y ==> 2(5-4y) + 13y = 15, after substitution.
10 - 8y + 13y = 15
==> 10 + 5y = 15 ==> 5y = 5 ==> y = 1 ==> x = 5 - 4*1 = 5 - 4 = 1.
Hence the solution is (1, 1).